Your backyard is 16 feet longer than it is wide. It's area is 4977 square feet. How wide is it?
Answer by Maths68(1474)

Area of the rectagle = width * length
Given
A=4977 squre feet
Width = W = L + 16feet
Length=L=?
A = W * L
4977 = (L + 16)*L
4977=L^2+16L
L^2+16L-4977=0
L^2+79L-63L-4977=0
L^2+79L-63L-4977=0
L(L+79)-63(L+79)=0
(L-63)(L+79)=0
L-63=0 or L+79=0
L=63 or L=-79 (impossible)
so length is 63feet
Given
W = L + 16feet
W=63+16
W=79feet